# Using the antenna impedance to estimate soil electrical parameters for the MIST global 21-cm experiment

### v1

@@ 3.1 Soil model
The soil was simulated as homogeneous, with a perfectly flat surface, and infinite in the horizontal directions and in depth. The electrical parameters of the soil model are its conductivity ($\sigma$) and relative permittivity ($\epsilon_{r}$). This model is a first-order approximation to the real soils at our sites, in which the surface is not perfectly flat and the properties, in general, vary as a function of space. The electrical properties of real soils, in general, also vary as a function of frequency (e.g., Smith-Rose & Appleton 1933; Smith-Rose 1935; Hipp 1974; Hoekstra & Delaney 1974; Hallikainen et al. 1985; Dobson et al. 1985; Francisca & Rinaldi 2003; Bobrov et al. 2015; Sutinjo et al. 2015). In this study, $\sigma$ and $\epsilon_{r}$ were modeled as constant in the range 25252525125125125125 MHz. We made this choice under the assumption that any variations of $\sigma$ and $\epsilon_{r}$ with frequency will have a small effect on the impedance compared to our experimental uncertainties. In Feko, the effect of the soil was incorporated analytically using Sommerfeld integrals as Green’s functions for solving the boundary conditions (Davidson 2011; Mosig & Michalski 2021).
3.2 Simulation settings
The simulations were done in the range 25252525125125125125 MHz to match the impedance measurements. However, carrying out the simulations at the $1$-MHz resolution of the measurements leads to computation times of $\sim 40$ minutes per simulation on our 64-core AMD Ryzen Threadripper PRO 5995WX computer. These computation times are impractical in a pipeline that requires a rigorous sampling of the soil parameter space. The computation time roughly scales with number of frequency points requested. To decrease the computation time, we reduced the resolution of the simulations to $10$ MHz and, thus, the number of points requested from $101$ to $11$. Using $11$ complex data points evenly spaced across 25252525125125125125 MHz effectively captures the main features of the antenna impedance and is sufficient for our analysis. Other settings used in the simulations include a fine automatic mesh size and double precision calculations. With this configuration, a simulation for a single combination of soil parameters was computed in $3.5$ minutes.
@@ 4 Parameter estimation
The conductivity and relative permittivity are estimated using the simplicial homology global optimization (SHGO) algorithm available in the SciPy Python library (Virtanen et al. 2020). This algorithm, introduced in Endres et al. (2018), enables derivative-free, constrained and unconstrained optimization. SHGO minimizes the cost function using global and local searches. For the global search, SHGO incorporates the simplicial (Hatcher 2002), Sobol (Sobol 1967; Joe & Kuo 2008), and Halton (Halton 1960) sampling methods, and can also accept externally provided sampling points. In this work we use the Sobol and Halton low-discrepancy sequences for the global search. These sequences are deterministic sets of sampling points with mathematical properties comparable to, and in some cases better than, the points from the popular Latin hypercube sampling McKay et al. (1979); Renardy et al. (2021). For the local search, we use sequential least squares quadratic programming (Kraft 1988; Nocedal & Wright 2006), which is the default method in SHGO.
As described in Sec. 2.4, in this initial implementation of the technique we do not know the dominant measurement uncertainties. In the absence of uncertainty information, with SHGO we minimize the sum of squared deviations (SSD),
@@
where $d$ represents the data, $m$ is the impedance model from a Feko simulation, and $i$ is the index of the $N$ points available across the frequency range. At each of the $11$ frequency channels where data and model are available, the resistance and reactance are treated as two independent points and, therefore, $N=22$. The best-fit soil parameters, denoted as $\hat{\sigma}$ and $\hat{\epsilon_{r}}$, are those that produce the lowest SSD, denoted as as SSD.
For each test site and sampling sequence, we conduct a parameter estimation with SHGO in two steps. In the initial step we run SHGO with wide allowed ranges for the parameters. Specifically, we use 001111 Sm-1 for $\sigma$ and 111120202020 for $\epsilon_{r}$. These ranges include the values from all the geophysical materials that could realistically be present in the soil at our test sites (Reynolds 2011). In the final step, we run SHGO again after constraining the parameter ranges around the best-fit values found in the initial step. In both steps, we use the default settings that define the stopping criterion in SHGO. These settings lead to a number of evaluations, $N_{\mathrm{ev}}$, under $250$ in each step. Our two-step fitting approach was chosen because, starting from a wide parameter space, it can yield best-fit parameters with relatively high numerical accuracy using a total $N_{\mathrm{ev}}<500$. If we instead use a single run to explore the wide parameter space, achieving a best-fit accuracy comparable to our two-step approach would require increasing $N_{\mathrm{ev}}$ potentially by orders of magnitude.
Our fitting approach with SHGO does not directly provide uncertainties in the best-fit parameters. We determine parameter uncertainties from the SHGO samples following a separate procedure. This procedure is described in Sec. 5.3.
Table 2: Results of the parameter estimation with SHGO.
Initial step Final step
Site Sampling $\hat{\sigma}$ [Sm-1] $\hat{\epsilon}_{r}$ SSD [$\Omega^{2}$] RMSD [$\Omega$] $N_{\mathrm{ev}}$ $\hat{\sigma}$ [Sm-1] $\hat{\epsilon}_{r}$ SSD [$\Omega^{2}$] RMSD [$\Omega$] $N_{\mathrm{ev}}$
A Sobol $10^{-5}$ $1$ $1132.0$ $7.2$ $131$ $0.0070$ $1.63$ $513.3$ $4.8$ $183$
A Halton $0.0054$ $6.95$ $1880.6$ $9.2$ $103$ $0.0064$ $1.82$ $531.8$ $4.9$ $151$
B Sobol $0.0234$ $8.57$ $1869.1$ $9.2$ $131$ $0.0144$ $9.61$ $1822.6$ $9.1$ $220$
B Halton $0.0210$ $8.12$ $1861.1$ $9.2$ $103$ $0.0165$ $9.77$ $1815.5$ $9.1$ $148$
C Sobol $0.0547$ $18.66$ $930.0$ $6.5$ $137$ $0.0494$ $12.73$ $825.9$ $6.1$ $188$
C Halton $0.0523$ $16.33$ $864.2$ $6.3$ $106$ $0.0479$ $14.69$ $834.1$ $6.2$ $148$
Figure 5: Samples used in the initial step of the parameter estimation. The colors of the samples represent the SSD. The three columns correspond to the three test sites. The top (bottom) row corresponds to the Sobol (Halton) samples. In each panel, the sample with the lowest SSD, denoted as SSD, is highlighted using a red circle. In all cases, SSD occurs for $\sigma<0.06$ Sm-1.
Figure 6: Same as Fig. 5 but for the final sampling step. This step enables the identification of the best-fit values with higher accuracy than the initial step. The color scale represents different SSD ranges for the three sites.
@@ 5 Results
5.1 Best-fit soil parameters
The results of the parameter estimation are summarized in Table 2. The table lists the best-fit soil parameters, $\hat{\sigma}$ and $\hat{\epsilon_{r}}$, and the corresponding SSD. In addition, the table shows the root mean square deviation (RMSD) between the best-fit impedance model and the data, $\mathrm{RMSD}^{\star}=\sqrt{\mathrm{SSD}^{\star}/N}$. Finally, the table shows $N_{\mathrm{ev}}$. These results are presented for the initial and final sampling steps, and for the Sobol and Halton samples.
Figure 5 shows the samples and SSDs for the initial step. The SSDs go from under 2,000 $\Omega^{2}$ to over 7,000 $\Omega^{2}$. For the three sites, the SSD decreases toward low $\sigma$, while across $\epsilon_{r}$ it remains more stable. The samples that produce the SSD are indicated using red circles. For Site A, as shown in Table 2, the best-fit values found in the initial step are $(\hat{\sigma},\hat{\epsilon_{r}})=(10^{-5}\;\mathrm{Sm}^{-1},1)$ for Sobol and $(0.0054\;\mathrm{Sm}^{-1},6.95)$ for Halton. These values are used as reference to determine adequate parameter ranges for the final step. The ranges selected for the final step are 000.020.020.020.02 Sm-1 for $\sigma$ and 111110101010 for $\epsilon_{r}$, which contain the initial best-fit values. For Sites B and C we follow a similar approach but decide to only restrict the $\sigma$ range. For these sites, the best-fit values found for $\epsilon_{r}$ in the initial step are higher than for Site A and the variation of the SSD across $\epsilon_{r}$ is low. We thus prefer to maintain 111120202020 as the range for $\epsilon_{r}$ in the final step. The $\sigma$ range is restricted to 000.040.040.040.04 Sm-1 for Site B and 000.080.080.080.08 Sm-1 for Site C.
Figure 6 shows the samples and SSDs for the final step. The final parameter spaces successfully capture the regions of lowest SSD and enable a denser sampling of these regions. The denser sampling leads to a higher accuracy in the best-fit values, which is reflected in the agreement at the $\sim 15\%$ level or better between the Sobol and Halton best-fit values. We report as the final best-fit values those that produce the lowest SSD considering both the Sobol and Halton sample sets. For Sites A, B, and C, the final best-fit values are $(\hat{\sigma},\hat{\epsilon}_{r})=(0.0070\;\mathrm{Sm}^{-1},1.63)$, $(0.0165\;\mathrm{Sm}^{-1},9.77)$, and $(0.0494\;\mathrm{Sm}^{-1},12.73)$, respectively. The best-fit values are shown in Table 3 including uncertainty ranges, which are estimated in Sec. 5.3.
5.2 Best-fit impedance simulations
@@
Figure 7 compares the measured impedances with the best-fit simulations. The data and simulations are shown at the common $10$-MHz resolution. The simulations show good general agreement with the data, with residuals that are within ±$\pm 25$ $\Omega$ and, for most of the points, within ±$\pm 10$ $\Omega$, especially for Sites A and C. The RMSD, which considers both the resistance and reactance, is $4.8$, $9.1$, and $6.1$ $\Omega$ for Sites A, B, and C, respectively. These values represent a better than $10\%$ agreement between best-fit simulations and data when using as a reference the peak absolute values of the resistance and reactance across our frequency range, which are $>100$ $\Omega$ for the three sites.
5.3 Uncertainties in the best-fit soil parameters
We determine uncertainties in the best-fit soil parameters reported in Sec. 5.1 using a Bayesian formalism. Bayes’ theorem indicates that the joint posterior probability density function (PDF) of the parameters given the data is
where $\mathcal{L}(d|\sigma,\epsilon_{r})$ represents the likelihood of the data given the parameters and $\pi(\sigma,\epsilon_{r})$ represents the prior PDF. We determine the parameter uncertainties taking into account our belief that $100\%$ of the probability volume is contained within the parameter space selected for the final step of the SHGO sampling, which is shown in Fig. 6. This belief is incorporated as a uniform prior across the final parameter space.
Computing $\mathcal{L}$ typically requires knowledge of the data uncertainty. In this paper we do not know the data uncertainty and compute $\mathcal{L}$ by making some simplifying assumptions. We first assume that the uncertainty across data points is uncorrelated and for each point follows a Gaussian distribution, which leads to the following likelihood:
$\mathcal{L}\propto\exp\left[-\frac{1}{2}\sum_{i=1}^{N}{\left(\frac{d_{i}-m_{i}}{\delta_{i}}\right)^{2}}\right],$ (4)
where $\delta$ is the standard deviation of the data uncertainty. We further assume that $\delta_{i}$ is the same across data points and equal to the RMSD reported in Sec. 5.2. We make this assumption because the RMSD is the only quantitative information we have about the discrepancy between the data and a good model for the data. With this additional assumption, the likelihood becomes
$\mathcal{L}\propto\exp\left[-\frac{\mathrm{SSD}}{2(\mathrm{RMSD}^{\star})^{2}}\right],$ (5)
where $\mathcal{L}$ depends on the soil parameters only through the SSD. Equation 5 enables us to map $\mathcal{L}$, and therefore the posterior PDFs, using the same samples and SSDs produced during the SHGO parameter fitting.
The posterior PDFs of the soil parameters are shown in Fig. 8. Panels (b), (e), and (h) show the joint PDFs for Sites A, B, and C, respectively. The joint PDFs are initially sampled using the existing Sobol and Halton samples, which are shown in the panels. Joint $68\%$ probability limits were computed after interpolating the joint PDFs at high resolution between the original samples. These limits are shown using the dashed cyan lines. In Fig. 8, the other panels show the marginalized PDFs for each individual parameter. For example, panels (a) and (c) show, respectively, the marginalized PDFs for $\sigma$ and $\epsilon_{r}$ at Site A. The marginalized PDFs were computed from the high-resolution interpolation of the joint PDFs. The $68\%$ uncertainties in the soil parameters, $\delta_{\sigma}$ and $\delta_{\epsilon_{r}}$, are determined from the marginalized PDFs; specifically, from the distance between the $68\%$ limits (dashed cyan lines) and the maximum a posteriori (MAP, dashed black lines).
Our results are summarized in Table 3. For all parameters and sites, the “+” and “-” $68\%$ ranges from the marginalized PDFs are equal to within $10\%$, and we therefore report as the $68\%$ uncertainty the average of the two. The table also reports the percent precision, which for $\hat{\sigma}$ is computed as $100\%\times(\delta_{\sigma}/\hat{\sigma})$. For Site B, the percent precision of $\hat{\sigma}$ and $\hat{\epsilon}_{r}$ is, respectively, $59\%$ and $32\%$. For Sites A and C, the percent precision of $\hat{\sigma}$ and $\hat{\epsilon}_{r}$ is, respectively, $20\%$ and $25\%$.
@@ 5.3 Uncertainties in the best-fit soil parameters
aBest-fit obtained as described in Sec. 5.1.

### v2

@@ 3.1 Soil Model
The soil was simulated as homogeneous, with a perfectly flat surface, and infinite in the horizontal directions and in depth. The electrical parameters of the soil model are its conductivity ($\sigma$) and relative permittivity ($\epsilon_{r}$) [^3] These parameters are components of the complex permittivity of a medium, $\epsilon$, defined as $\epsilon=\epsilon^{\prime}-j\epsilon^{\prime\prime}=\epsilon_{0}\epsilon_{r}-j\frac{\sigma}{2\pi f}$, where $\epsilon_{0}$ is the permittivity of vacuum and $f$ is frequency (von Hippel, 1995; Pozar, 2005).. This soil model is a first-order approximation to the real soils at our sites, in which the surface is not perfectly flat and the properties, in general, vary as a function of space. The electrical properties of real soils, in general, also vary as a function of frequency (e.g., Smith-Rose & Appleton, 1933; Smith-Rose, 1935; Hipp, 1974; Hoekstra & Delaney, 1974; Hallikainen et al., 1985; Dobson et al., 1985; Francisca & Rinaldi, 2003; Bobrov et al., 2015; Sutinjo et al., 2015) and time (Pattison et al., 2025). In this study, $\sigma$ and $\epsilon_{r}$ were modeled as constant in the range 2525125125 MHz. We made this choice under the assumption that any variations of $\sigma$ and $\epsilon_{r}$ with frequency will have a small effect on the impedance compared to our experimental uncertainties. In Feko, the effect of the soil was incorporated analytically using Sommerfeld integrals as Green’s functions for solving the boundary conditions (Davidson, 2011; Mosig & Michalski, 2021).
Figure 5: Feko simulations conducted to study the sensitivity of $\Gamma$ and $Z$ to the soil parameter values. The reference case (black line) corresponds to the geometry simulated in free space. In the reference case, the structure observed across frequency with a period of $\sim$$7.5$ MHz is produced by reflections from the end of the $20$ m cable. This structure is suppressed when $\sigma$ or $\epsilon_{r}$ are increased because reflections from the soil, which span a wide range of delays, have a smoothing effect and reduce the impact from the cable alone.
3.2 Sensitivity to the Soil Parameters
To develop intuition about the sensitivity of the antenna reflection coefficient and impedance to the soil parameters, in Figure 5 we show Feko simulations for five combinations of soil parameter values. The reference parameter combination is $(\sigma,{\epsilon}_{r})=(0\;\mathrm{Sm}^{-1},1)$, which is equivalent to free space. Relative to the reference, in two cases we increase $\sigma$ and in the other two we increase $\epsilon_{r}$. The features across frequency that are common among the simulations are those also identified in the measurements (Section 2.4); i.e., a dip in $|\Gamma|$, a fast transition in $\angle\Gamma$, a peak in $R$, and a fast transition in $X$. We note that these characteristic features are present even when the antenna is in free space, and can therefore be regarded as intrinsic to the antenna structure instead of imposed by the soil. An additional noteworthy feature observed in the free space simulation is the repetitive structure seen across frequency in $|\Gamma|$, $\angle\Gamma$, $R$, and $X$. This structure has a period of $\sim$$7.5$ MHz, which in free space corresponds to a wavelength or travel distance of $\sim$$40$ m. The structure is therefore attributed to reflections from the end of the $20$ m cable included in the full geometry. When $\sigma$ or $\epsilon_{r}$ increase compared to free space, the periodic structure is suppressed. This behavior occurs because the soil surface also produces reflections, which span a wide range of delays and have a smoothing effect across frequency. Figure 5 also shows that increasing $\sigma$ or $\epsilon_{r}$ produces a higher $|\Gamma|$, a smoother transition in $\angle\Gamma$, a narrower and higher peak in $R$, and a larger amplitude in the transition of $X$. Finally, we note the following patterns in the impedance. Increasing $\sigma$ ($\epsilon_{r}$) shifts the peak of $R$ to lower (higher) frequencies. Regarding $X$, increasing $\sigma$ shifts the high-frequency knee of the transition (dip at $\sim$$80$ MHz) to lower frequencies. Conversely, increasing $\epsilon_{r}$ shifts the low-frequency knee of the transition (peak at $\sim$60 MHz) to higher frequencies.
@@ 4 Parameter Estimation
We estimate the soil conductivity and relative permittivity at our test sites by fitting the measured antenna impedances with models from Feko simulations. As described in Section 2.4, in this initial implementation of the technique we do not know the dominant measurement uncertainties. For this reason, instead of minimizing chi squared, in the fits we minimize the sum of squared deviations (SSD),
@@
where $d$ represents the data, $m$ is the impedance model from a Feko simulation, and $i$ is the index of the $N$ points available across the frequency range. The best-fit soil parameters, which we call $\hat{\sigma}$ and $\hat{\epsilon}_{r}$, are those that produce the lowest SSD, denoted by SSD.
Carrying out Feko simulations in the range 2525125125 MHz at the $1$ MHz resolution of the measurements leads to computation times of $\sim$$40$ minutes per simulation on our 64-core AMD Ryzen Threadripper PRO 5995WX computer. These computation times are impractical in a pipeline requiring a rigorous sampling of the soil parameter space. The computation time roughly scales with number of frequency points requested. To decrease the computation time, we reduced the resolution of the simulations to $10$ MHz and, thus, the number of points requested from $101$ to $11$. Using $11$ complex data points evenly spaced across 2525125125 MHz effectively captures the main features of the antenna impedance and is sufficient for our analysis. With this configuration, a simulation for a single combination of soil parameters is computed in $3.5$ minutes. At each of the $11$ frequency channels where data and model are available, the resistance and reactance are treated as two independent points and, therefore, $N=22$.
Since we do not know the measurement uncertainties when conducting the fits, we cannot compute the data likelihood and posterior probability density function (PDF) of $\sigma$ and $\epsilon_{r}$ with high fidelity. Therefore, it is not justified to fit the soil parameters using techniques such as nested sampling or MCMC, which are designed to map the PDFs and would likely require thousands of computationally expensive Feko simulations. We instead use a method that conducts a global search to determine the best-fit values and requires only a few hundred simulations. This method is described in Appendix B. At the core of the method is the simplicial homology global optimization (SHGO) algorithm. In this method, we initially fit the soil parameters considering a wide parameter space: 011 Sm-1 for $\sigma$ and 112020 for $\epsilon_{r}$. This space includes the values from all the geophysical materials that could realistically be present in the soil at our three test sites (Reynolds, 2011). This initial step identifies the general regions where the best-fits lie but does not provide accurate estimates. To increase the fitting accuracy, we carry out a second fitting step where the global search is conducted within a smaller parameter space. This smaller space is chosen based on the results from the first step. SHGO does not directly provide uncertainties in the best-fit parameters. However, we determine first-order parameter uncertainties in a separate procedure using the samples acquired during the fits. This procedure is described in Section. 5.2.
@@ 5 Results
5.1 Best-fit soil parameters and impedance simulations
For Sites A, B, and C, the best-fit parameter values found with SHGO are $(\hat{\sigma},\hat{\epsilon}_{r})=(0.0070\;\mathrm{Sm}^{-1},1.63)$, $(0.0165\;\mathrm{Sm}^{-1},9.77)$, and $(0.0494\;\mathrm{Sm}^{-1},12.73)$, respectively. The corresponding $\mathrm{SSD}^{\star}$ are $513.3$, $1815.5$, and $825.9$ $\Omega^{2}$. The total number of Feko simulations conducted to determine the best-fit values at each of the three test sites was $\leq 602$. More details of the fits are provided in Appendix B. The best-fit parameter values are shown in Table 2 including uncertainty ranges, which are computed as described in Sec. 5.2.
@@ 5.1 Best-fit soil parameters and impedance simulations
Figure 6 compares the measured impedances with the best-fit simulations. The data and simulations are shown at the common $10$ MHz resolution. The simulations show good general agreement with the data, with residuals that are within $\pm 25$ $\Omega$, and for most of the points within $\pm 10$ $\Omega$, especially for Sites A and C. We also quantify the residuals using the root mean square deviation (RMSD) between the best-fit models and the data; specifically, $\mathrm{RMSD}^{\star}=\sqrt{\mathrm{SSD}^{\star}/N}$. The RMSD considers both the resistance and reactance, and is $4.8$, $9.1$, and $6.1$ $\Omega$ for Sites A, B, and C, respectively. These values represent a better than $10\%$ agreement between best-fit simulations and data when using as a reference the peak absolute values of the resistance and reactance across our frequency range, which are >100>100 $\Omega$ for the three sites. When considering the origin of the residuals, we cannot rule out the possibility that at least part of them is due to the inadequacy of our homogeneous soil models to simulate the real soils. Unfortunately, our interpretations are fundamentally restricted by the lack of information about our measurement uncertainties. The quantification of these uncertainties should therefore be prioritized in future implementations of this technique.
@@
aBest-fit obtained directly with SHGO.
@@
5.2 Uncertainties in the best-fit soil parameters
We estimate first-order uncertainties in the best-fit soil parameters from estimates of the parameters’ posterior PDFs. Bayes’ theorem indicates that the joint posterior PDF of the parameters given the data is
where $\mathcal{L}(d|\sigma,\epsilon_{r})$ represents the likelihood of the data given the parameters and $\pi(\sigma,\epsilon_{r})$ represents the prior PDF. We compute the posterior PDFs using the samples produced during the SHGO parameter fitting. As a prior we use uniform PDFs that span the parameter spaces selected for the final fitting step with SHGO (Appendix B). Computing $\mathcal{L}$ requires knowledge of the measurement uncertainties. Since we do not have this information we make simplifying assumptions. Specifically, we first assume that the measurements are affected by errors that are random, Gaussian, and uncorrelated across frequency. We further assume that the errors have a frequency-independent standard deviation equal to the RMSD reported in Sec. 5.1. With these assumptions, the likelihood becomes
$\mathcal{L}\propto\exp\left[-\frac{\mathrm{SSD}}{2(\mathrm{RMSD}^{\star})^{2}}\right].$ (4)
The structure of the residuals in Figure 6 does not appear random, Gaussian, or uncorrelated across frequency, but we make these assumptions about the errors as an approximation in the absence of independently determined information. We caution, therefore, that the parameter uncertainties determined here should only be regarded as first-order estimates.
The posterior PDFs computed for the soil parameters are shown in Figure 7. Panels (b), (e), and (h) show the joint PDFs for Sites A, B, and C, respectively. The samples shown in the panels come from the SHGO fits. The PDF surfaces revealed by the samples are smooth and contain compact regions of high probability that resemble Gaussians. Joint $68\%$ probability limits were computed after interpolating the joint PDFs at high resolution between the SHGO samples. These limits are shown using the dashed cyan lines. In Figure 7, the other panels show marginalized PDFs for each individual parameter. For example, panels (a) and (c) show, respectively, the marginalized PDFs for $\sigma$ and $\epsilon_{r}$ at Site A. The marginalized PDFs were computed from the high-resolution interpolation of the joint PDFs. The $68\%$ uncertainties in the soil parameters, denoted by $\delta_{\sigma}$ and $\delta_{\epsilon_{r}}$, are determined from the marginalized PDFs; specifically, from the distance between the $68\%$ limits (dashed cyan lines) and the maximum a posteriori (MAP, dashed black lines).
[…]
5.3 Comparison with expectations
In this initial effort to demonstrate the antenna impedance technique we do not attempt to offer a geophysical interpretation of our results shown in Table 2. We note, however, that our results are in broad agreement with expectations from the soil composition at the test sites. Sand and gravel, found at the surface of Sites A and B, have reported conductivity $\lesssim 0.01$ Sm-1 and relative permittivity in the range $\sim$22-$30$. For clay, found at the surface of Site C, the conductivity is in the range $0^{-4}11$ Sm-1 and the relative permittivity in the range $\sim$224040. Within these ranges, the specific electrical parameter values for soils composed of sand, gravel, or clay are strongly determined by the soil moisture (Reynolds, 2011).
Our conductivity estimate for Site A is not expected to match very closely the results from Alegría Calvo et al. (2013). First, being $650$ m apart, the two spots do not necessarily have similar soil characteristics. Second, while our result is valid for the range 2525125125 MHz, in ERT the resistivity —inverse of conductivity— is estimated in the DC regime [^4] Electrodes are used to inject direct currents into the soil and measure the resulting voltages between different spots (Slichter, 1933; Koh et al., 1997; IEEE, 2011; Herring et al., 2023). Sometimes, the polarization of the injected current is reversed at frequencies $\lesssim 100$ Hz to reduce the effect of electrolytic polarization (build up of anions and cations around the electrodes) and telluric currents (currents in the ground mainly produced by the presence and fluctuation of the Earth’s magnetosphere) (IEEE, 2011; Reynolds, 2011; Clement et al., 2020).. Despite not expecting a good agreement, such an agreement would be mathematically possible when accounting for the spatial dependence of the results. Our result represents a weighted average of the soil conductivity over space, where the weights corresponds to the sensitivity of the antenna at different coordinates in the soil. Reporting the pattern for the antenna sensitivity in the soil falls outside the scope of this paper. However, using a reasonable set of weights for the conductivities of the four layers of the ERT model it would be possible to obtain a weighted average close to our value. A realistic situation consists of the antenna having higher sensitivity toward the top of the soil. Therefore, assigning as an example weights of $96\%$ to the top layer, $4\%$ to the second layer, and $0\%$ to the other layers, produces a weighted average of $0.0076$ Sm-1, consistent with our result $0.0070\pm 0.0014$ Sm-1.